OpenStudy (anonymous):

When the equation of the line tangent to the curve y=(kx+8)/(k+x) at x=-2 is y=x+4. What is the value of k? A. -3 B. -1 C. 1 D. 3 E. 4

OpenStudy (anonymous):

\[y=\frac{kx+8}{k+x}\to f^{\prime}(x) = \frac{(k)(k+x)-(1)(kx+8)}{(k+x)^2}\] So when \[f^{\prime}(-2)=1=\frac{(k)(k+(-2))-(1)(k(-2)+8)}{(k+(-2))^2}\] solve for k.

OpenStudy (anonymous):

f'(-2)=(k^2-8)/(k-2)^2 what am i supposed to do next?

OpenStudy (tkhunny):

At x = -2, the value of the derivative is +1. Why do I know that?

OpenStudy (anonymous):

Um...I have no idea...

OpenStudy (tkhunny):

It's hiding in the problem statement: " x=-2 is y=x+4" What's the slope of that line that was given as reference? There it is!

OpenStudy (anonymous):

can you simplify the derivative dy/dx = ??? what u got?

OpenStudy (anonymous):

i got k=3?

OpenStudy (anonymous):

yep...

OpenStudy (tkhunny):

Did you circle back around and prove it in the original equation? That will build your confidence.

OpenStudy (anonymous):

I realized that the derivative has to equal to the slope of the tangent line which is 1.

OpenStudy (tkhunny):

Victory Prize! Go challenge the next one, now.

OpenStudy (anonymous):

thank you!